# A Calculus of Inheritance

### 10 Mar 2026 v6

@@ B.6. Adequacy
Base case ($k=0$): $M$ is already an abstraction $\lambda x.\,M^{\prime}$. Then $\mathcal{T}(M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$, whose root has $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$. These labels are directly in $\mathtt{defines}$ of the root, so the recursive evaluation returns them without further recursion, and $\mathcal{T}(M)$ converges at depth $n=0$.
Inductive step ($k\geq 1$): $M$ is not an abstraction, so $M$ is either a tail call, an application $\mathbf{let}$-binding, or a value $\mathbf{let}$-binding. Each form corresponds to one head $\beta$-step in the underlying $\lambda$-term. Let $N$ be the head reduct, so $M\to_{h}N\to^{k-1}_{h}\lambda x.\,M^{\prime}$. Note that $N$ is closed: head reduction substitutes a closed value into a term whose only free variable is the substitution target, preserving closedness. By Lemma B.3 (tail-call, application let-binding, or value let-binding variant), the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$. By the induction hypothesis, $\mathcal{T}(N)$ converges at depth $\leq k-1$, so $\mathcal{T}(M)$ converges at depth $\leq k$.
@@
Base case ($n=0$): the root of $\mathcal{T}(M)$ has abstraction shape $\{\mathrm{argument},\mathrm{result}\}$. Only the abstraction rule $\mathcal{T}(\lambda x.\,M^{\prime})$ produces a root with own property $\mathrm{argument}$. (For tail calls and $\mathbf{let}$-bindings, the root has $\mathtt{defines}=\{\mathrm{tailCall},\mathrm{result}\}$ or $\{x,\mathrm{result}\}$, and $\mathrm{argument}$ cannot appear via $\mathtt{supers}$ because $\mathtt{inherits}=\varnothing$ at the root, giving $\mathtt{bases}(\text{root})=\varnothing$.) Therefore $M$ is an abstraction, hence already in head normal form.
Inductive step ($n\geq 1$): the abstraction shape appears at depth $n$ but not at depth 0. By the base case argument, $M$ is not an abstraction, so at the head position of this closed top-level ANF term, $M$ is either a tail call, an application $\mathbf{let}$-binding, or a value $\mathbf{let}$-binding. In each case, the head form is a $\beta$-redex in the underlying $\lambda$-term: the function position is occupied by a closed ANF value, and the only closed ANF value that can appear at head position is a $\lambda$-abstraction. By Lemma B.3 (tail-call, application let-binding, or value let-binding variant), the properties at depth n1n-1 in $\mathcal{T}(N)$ match those at depth $n$ in $\mathcal{T}(M)$, where $N$ is the head reduct. So $\mathcal{T}(N)$ converges at depth n1n-1. Since $M$ is closed and head reduction preserves closedness (substituting a closed value for the only free variable), $N$ is a closed ANF term. By the induction hypothesis, $N$ has a head normal form, and since $M\to_{h}N$, so does $M$. ∎
@@ B.7. First-Order Semantics of the Lazy $\lambda$-Calculus
The subscript $\lambda$ emphasizes that the quantification ranges over $\lambda$-calculus contexts, not over all inheritance-calculus contexts. This definition observes only convergence (a single bit), but quantifying over all $\lambda$-contexts makes it a fine-grained equivalence: different contexts can supply different arguments and project different results, probing every aspect of the term’s behavior within the $\lambda$-definable fragment.
Proof of Theorem 4.3.
@@
($\Leftarrow$) Suppose $\mathrm{BT}(M)=\mathrm{BT}(N)$. Böhm tree equivalence is a congruence (Barendregt, 1984), so for every context $C[\cdot]$, $\mathrm{BT}(C[M])=\mathrm{BT}(C[N])$. In particular, $C[M]$ has a head normal form iff $C[N]$ does. By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ iff $\mathcal{T}(C[N]){\Downarrow}$. Hence $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$.
($\Rightarrow$) Suppose $\mathrm{BT}(M)\neq\mathrm{BT}(N)$. Böhm tree equivalence coincides with observational equivalence for the lazy $\lambda$-calculus (Abramsky and Ong, 1993): there exists a context $C[\cdot]$ such that $C[M]$ converges and $C[N]$ diverges (or vice versa). By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ and $\mathcal{T}(C[N]){\Uparrow}$. Hence $\mathcal{T}(M)\not\approx_{\lambda}\mathcal{T}(N)$. ∎
The proof rests on two pillars: Adequacy (Theorem 4.2), which is our contribution, and the classical result that Böhm tree equivalence equals observational equivalence for the lazy $\lambda$-calculus (Abramsky and Ong, 1993). The prior higher-order frameworks are thereby shown to be faithful to the first-order semantics, without needing to define or compare absolute paths inside mixin trees.
@@ Appendix C Expressive Asymmetry: Proofs
Let $\mathscr{U}$ be the language whose constructors are the union of the constructors of inheritance-calculus, the constructors of the lazy $\lambda$-calculus ($\lambda$-abstraction, application, variable), and $\mathbf{let}$-binding. The semantics of $\mathscr{U}$ is given by translating every $\mathscr{U}$-program to inheritance-calculus in three steps:
@@
    Convert every maximal $\lambda$-calculus subexpression (i.e. every subtree whose root and all internal nodes use only $\lambda$-calculus constructors) to A-normal form (Flanagan et al., 1993).
@@
A $\mathscr{U}$-program $P$ converges, $\mathrm{eval}_{\mathscr{U}}(P)$, iff its translation converges. The lazy $\lambda$-calculus and inheritance-calculus are conservative restrictions of $\mathscr{U}$: each is obtained by removing the other’s constructors. For a pure inheritance-calculus program, steps (1) and (2) are vacuous, so $\mathrm{eval}_{\mathscr{U}}$ agrees with the inheritance-calculus semantics. For a pure $\lambda$-calculus program, the composition of steps (1)–(3) agrees with lazy evaluation by Adequacy (Theorem 4.2).
@@
    $\lambda$-bound variable ($x$ with de Bruijn index $n^{\prime}$). This is a nullary construct (no subexpressions). The syntactic abstraction is the constant $A=\uparrow^{\!n^{\prime}}.\mathrm{argument}$.
@@
    $\lambda$-abstraction ($\lambda x.\,M$). Unary in the body $M$. $A(\cdot_{1})=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\cdot_{1}\}$.
@@
    Semantics preserved: For pure $\lambda$-terms, steps (1) and (2) of Definition C.3 are vacuous and step (3) yields lazy evaluation by Adequacy (Theorem 4.2), so $\mathrm{eval}_{\mathscr{U}}$ agrees with $\mathrm{eval}_{L_{0}}$ on $L_{0}$-phrases.
@@
By Böhm tree equivalence (Abramsky, 1990) (Theorem 4.3), for closed $\lambda$-terms $M$ and $N$: $M\cong_{0}N$ iff $\mathrm{BT}(M)=\mathrm{BT}(N)$.
@@
Since $\mathrm{eqFalse}$ (respectively $\mathrm{eqTrue}$) inherits $\mathrm{eq}.\mathrm{result}$, and $\mathrm{tailCall}$ inherits $\mathrm{eqFalse}.\mathrm{result}$ (the $\lambda b$ scope), the $\mathrm{repr}$ record placed by the mixin at the $\lambda b$ scope level propagates through the inheritance chain to $\mathrm{tailCall}$. In $C[M]$: the $\mathrm{eqFalse}$ binding supplies $\mathrm{argument}\mapsto f_{1}.\mathrm{result}$ (i.e. $\mathrm{false}$) to the $\lambda a$ scope, so $\mathrm{firstOperand}$ evaluates to $\mathrm{false}=\lambda t.\,\lambda f.\,f$. In $C[N]$: the $\mathrm{eqTrue}$ binding supplies $\mathrm{argument}\mapsto\mathrm{true}$, so $\mathrm{firstOperand}$ evaluates to $\mathrm{true}=\lambda t.\,\lambda f.\,t$.
@@ Conclusion
Appendix D Multi-Path Self-Reference in Other Systems
This appendix demonstrates that two representative systems, Scala 3 and the NixOS module system, reject the multi-path self-reference pattern that inheritance-calculus handles naturally.
@@ Appendix D Multi-Path Self-Reference in Other Systems
The rejection is not a surface-level restriction but a consequence of DOT’s type-theoretic foundations (Amin et al., 2016). In DOT, an object is constructed as $\nu(x:T)\,d$, where the self variable $x$ binds to a single object. Object1.MyInner and Object2.MyInner are distinct path-dependent types, each constraining the outer self-reference to a different object. Their intersection requires outer to simultaneously return Object1 and Object2, but DOT’s self variable is single-valued, making this intersection unrealizable.
@@
This is well-defined in inheritance-calculus. The reference $\mathrm{MyOuter}.\mathbf{this}$ inside $\mathrm{outer}$ has de Bruijn index $n=1$: starting from $\mathrm{outer}$’s enclosing scope $\mathrm{MyInner}$, one $\mathtt{this}$ step reaches $\mathrm{MyOuter}$. When $\mathrm{HasMultipleOuters}$ inherits from both $\mathrm{Object1}.\mathrm{MyInner}$ and $\mathrm{Object2}.\mathrm{MyInner}$, the $\mathtt{this}$ function (equation 6) resolves $\mathrm{MyOuter}.\mathbf{this}$ by searching through $\mathtt{supers}(\mathrm{HasMultipleOuters})$ for override paths matching $\mathrm{MyInner}$’s definition site. It finds two inheritance-site paths, one through $\mathrm{Object1}$ and one through $\mathrm{Object2}$, and returns both. Both paths lead to records that inherit from the same $\mathrm{MyOuter}$, so querying $\mathrm{HasMultipleOuters}.\mathrm{outer}$ yields the $\mathtt{properties}$ of $\mathrm{MyOuter}$ regardless of which route is taken. Because attribute resolution aggregates results via set union (which is intrinsically idempotent), identical property definitions from the two routes merge without conflict. Thus, the calculus eliminates the need for any artificial route-priority or linearization rules.
@@
The rejection occurs because Object1.MyInner and Object2.MyInner each carry the full staticModules of the MyInner deferred module, including the declaration options.outer. When MyObjectA imports both, the module system’s mergeOptionDecls encounters two declarations of the same option and raises a static error; it cannot express that both declarations originate from the same trait and should be unified.
The rejection is analogous to Scala’s “conflicting base types” error above: both systems assume that each option or type member has a single declaration site, and reject the multi-path situation where two inheritance routes introduce the same declaration independently. In inheritance-calculus, the same pattern is well-defined because $\mathtt{overrides}$ (equation 3) recognizes that both routes lead to the same definition, and $\mathtt{supers}$ (equation 2) collects both inheritance-site contexts without duplication.
Appendix E First-Order Semantics of the ANF Lazy $\lambda$-Calculus
This appendix presents a first-order semantics of the lazy $\lambda$-calculus in A-normal form (ANF). The semantics is given by mutually recursive set-comprehension equations whose well-definedness follows from the same termination argument as in Appendix A. The central idea is that function application is modeled as tree grafting: applying a function record to an argument splices the function’s subtree into the call site, and variable references navigate the resulting tree structure.
ANF is essential. Traditional $\lambda$-calculus syntax allows anonymous subexpressions such as (fx)(f\;x) in (fx)y(f\;x)\;y that have no label and cannot form paths. ANF restores path-addressability by binding every intermediate result to a name, so that every node in the evaluation tree has a unique path address.

### 16 Mar 2026 v7

@@ B.6. Adequacy
Definition B.6 (Inheritance-convergence).
in the semantics of Section 3.
Theorem B.7 (Adequacy).
@@
Base case ($k=0$)
$M$ is already an abstraction $\lambda x.\,M^{\prime}$. Then $\mathcal{T}(M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$, whose root has $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$. These labels are directly in $\mathtt{defines}$ of the root, so the recursive evaluation returns them without further recursion, and $\mathcal{T}(M)$ converges at depth $n=0$.
Inductive step ($k\geq 1$)
$M$ is not an abstraction, so $M$ is either a tail call, an application $\mathbf{let}$-binding, or a value $\mathbf{let}$-binding. Each form corresponds to one head $\beta$-step in the underlying $\lambda$-term. Let $N$ be the head reduct, so $M\to_{h}N\to^{k-1}_{h}\lambda x.\,M^{\prime}$. Note that $N$ is closed: head reduction substitutes a closed value into a term whose only free variable is the substitution target, preserving closedness. By Lemma B.3 (tail-call, application let-binding, or value let-binding variant), the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$. By the induction hypothesis, $\mathcal{T}(N)$ converges at depth $\leq k-1$, so $\mathcal{T}(M)$ converges at depth $\leq k$.
@@
Base case ($n=0$)
The root of $\mathcal{T}(M)$ has abstraction shape $\{\mathrm{argument},\mathrm{result}\}$. Only the abstraction rule $\mathcal{T}(\lambda x.\,M^{\prime})$ produces a root with own property $\mathrm{argument}$. (For tail calls and $\mathbf{let}$-bindings, the root has $\mathtt{defines}=\{\mathrm{tailCall},\mathrm{result}\}$ or $\{x,\mathrm{result}\}$, and $\mathrm{argument}$ cannot appear via $\mathtt{supers}$ because $\mathtt{inherits}=\varnothing$ at the root, giving $\mathtt{bases}(\text{root})=\varnothing$.) Therefore $M$ is an abstraction, hence already in head normal form.
Inductive step ($n\geq 1$)
The abstraction shape appears at depth $n$ but not at depth 0. By the base case argument, $M$ is not an abstraction, so at the head position of this closed top-level ANF term, $M$ is either a tail call, an application $\mathbf{let}$-binding, or a value $\mathbf{let}$-binding. In each case, the ANF grammar requires the function position to be a value $V_{1}$ ($V::=x\mid\lambda x.\,M$). Since $M$ is closed, $V_{1}$ cannot be a free variable, so $V_{1}$ must be a $\lambda$-abstraction; the head form is therefore a $\beta$-redex. By Lemma B.3 (tail-call, application let-binding, or value let-binding variant), the properties at depth n1n-1 in $\mathcal{T}(N)$ match those at depth $n$ in $\mathcal{T}(M)$, where $N$ is the head reduct. So $\mathcal{T}(N)$ converges at depth n1n-1. Since $M$ is closed and head reduction preserves closedness , which substitutes a closed value for the only free variable, $N$ is a closed ANF term. By the induction hypothesis, $N$ has a head normal form, and since $M\to_{h}N$, so does $M$. ∎
@@ B.7. First-Order Semantics of the Lazy $\lambda$-Calculus
The subscript $\lambda$ emphasizes that the quantification ranges over $\lambda$-calculus contexts, not over all inheritance-calculus contexts. This definition observes only convergence [^33] Convergence provides exactly a single bit of information. , but quantifying over all $\lambda$-contexts makes it a fine-grained equivalence: different contexts can supply different arguments and project different results, probing every aspect of the term’s behavior within the $\lambda$-definable fragment.
Theorem B.9 (First-order semantics of the lazy $\lambda$-calculus).
For closed $\lambda$-terms $M$ and $N$:
Proof of Theorem B.9.
@@
($\Leftarrow$) Suppose $\mathrm{BT}(M)=\mathrm{BT}(N)$. Böhm tree equivalence is a congruence (Barendregt, 1984), so for every context $C[\cdot]$, $\mathrm{BT}(C[M])=\mathrm{BT}(C[N])$. In particular, $C[M]$ has a head normal form iff $C[N]$ does. By Theorem B.7, $\mathcal{T}(C[M]){\Downarrow}$ iff $\mathcal{T}(C[N]){\Downarrow}$. Hence $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$.
($\Rightarrow$) Suppose $\mathrm{BT}(M)\neq\mathrm{BT}(N)$. Böhm tree equivalence coincides with observational equivalence for the lazy $\lambda$-calculus (Abramsky and Ong, 1993; Abramsky and McCusker, 1995): there exists a context $C[\cdot]$ such that $C[M]$ converges and $C[N]$ diverges, or vice versa. By Theorem B.7, $\mathcal{T}(C[M]){\Downarrow}$ and $\mathcal{T}(C[N]){\Uparrow}$. Hence $\mathcal{T}(M)\not\approx_{\lambda}\mathcal{T}(N)$. ∎
The proof rests on two pillars: Adequacy (Theorem B.7), which is our contribution, and the classical result that Böhm tree equivalence equals observational equivalence for the lazy $\lambda$-calculus (Abramsky and Ong, 1993; Abramsky and McCusker, 1995). Neither pillar requires defining or comparing absolute paths inside mixin trees.
@@ Appendix C Expressive Asymmetry: Proofs
Let $\mathscr{U}$ be the language whose constructors are the union of the constructors of inheritance-calculus, the constructors of the lazy $\lambda$-calculus , including $\lambda$-abstraction, application, and variable, and $\mathbf{let}$-binding. The semantics of $\mathscr{U}$ is given by translating every $\mathscr{U}$-program to inheritance-calculus in three steps:
@@
    Convert every maximal $\lambda$-calculus subexpression (that is, every subtree whose root and all internal nodes use only $\lambda$-calculus constructors) to A-normal form (Flanagan et al., 1993).
@@
A $\mathscr{U}$-program $P$ converges, $\mathrm{eval}_{\mathscr{U}}(P)$, iff its translation converges.
Remark C.4.
The lazy $\lambda$-calculus and inheritance-calculus are conservative restrictions of $\mathscr{U}$: each is obtained by removing the other’s constructors. For a pure inheritance-calculus program, steps (1) and (2) are vacuous, so $\mathrm{eval}_{\mathscr{U}}$ agrees with the inheritance-calculus semantics. For a pure $\lambda$-calculus program, the composition of steps (1)–(3) agrees with lazy evaluation by Adequacy (Theorem B.7).
@@
    $\lambda$-bound variable ($x$ with de Bruijn index $n^{\prime}$). This is a nullary construct and therefore has no subexpressions. The syntactic abstraction is the constant $A=\uparrow^{\!n^{\prime}}.\mathrm{argument}$.
@@
    $\lambda$-abstraction, denoted $\lambda x.\,M$. Unary in the body $M$. $A(\cdot_{1})=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\cdot_{1}\}$.
@@
    Semantics preserved: For pure $\lambda$-terms, steps (1) and (2) of Definition C.3 are vacuous and step (3) yields lazy evaluation by Adequacy (Theorem B.7), so $\mathrm{eval}_{\mathscr{U}}$ agrees with $\mathrm{eval}_{L_{0}}$ on $L_{0}$-phrases.
@@
By Böhm tree equivalence (Abramsky and Ong, 1993; Abramsky and McCusker, 1995) (Theorem B.9), for closed $\lambda$-terms $M$ and $N$: $M\cong_{0}N$ iff $\mathrm{BT}(M)=\mathrm{BT}(N)$.
@@
Since $\mathrm{eqFalse}$ (respectively $\mathrm{eqTrue}$) inherits $\mathrm{eq}.\mathrm{result}$, and $\mathrm{tailCall}$ inherits $\mathrm{eqFalse}.\mathrm{result}$ (the $\lambda b$ scope), the $\mathrm{repr}$ record placed by the mixin at the $\lambda b$ scope level propagates through the inheritance chain to $\mathrm{tailCall}$. In $C[M]$: the $\mathrm{eqFalse}$ binding supplies $\mathrm{argument}\mapsto f_{1}.\mathrm{result}$ (that is, $\mathrm{false}$) to the $\lambda a$ scope, so $\mathrm{firstOperand}$ evaluates to $\mathrm{false}=\lambda t.\,\lambda f.\,f$. In $C[N]$: the $\mathrm{eqTrue}$ binding supplies $\mathrm{argument}\mapsto\mathrm{true}$, so $\mathrm{firstOperand}$ evaluates to $\mathrm{true}=\lambda t.\,\lambda f.\,t$.
@@ Conclusion
Appendix D Multi-Target $\mathtt{this}$ Resolution in Other Systems
This appendix demonstrates that two representative systems, Scala 3 and the NixOS module system, reject the multi-target $\mathtt{this}$ resolution pattern that inheritance-calculus handles naturally.
@@ Appendix D Multi-Target $\mathtt{this}$ Resolution in Other Systems
The rejection is not a surface-level restriction but a consequence of DOT’s type-theoretic foundations (Amin et al., 2016). In DOT, an object is constructed as $\nu(x:T)\,d$, where the self variable $x$ binds to a single object. Object1.MyInner and Object2.MyInner are distinct path-dependent types, each constraining the outer $\mathtt{this}$ to a different object. Their intersection requires outer to simultaneously return Object1 and Object2, but DOT’s self variable is single-valued, making this intersection unrealizable.
@@
This is well-defined in inheritance-calculus. The reference $\mathrm{MyOuter}.\mathbf{this}$ inside $\mathrm{outer}$ has de Bruijn index $n=1$: starting from $\mathrm{outer}$’s enclosing scope $\mathrm{MyInner}$, one $\mathtt{this}$ step reaches $\mathrm{MyOuter}$. When $\mathrm{HasMultipleOuters}$ inherits from both $\mathrm{Object1}.\mathrm{MyInner}$ and $\mathrm{Object2}.\mathrm{MyInner}$, the $\mathtt{this}$ function (equation 6) resolves $\mathrm{MyOuter}.\mathbf{this}$ by searching through $\mathtt{supers}(\mathrm{HasMultipleOuters})$ for override paths matching $\mathrm{MyInner}$’s definition site. It finds two inheritance-site paths, one through $\mathrm{Object1}$ and one through $\mathrm{Object2}$, and returns both. Both paths lead to records that inherit from the same $\mathrm{MyOuter}$, so querying $\mathrm{HasMultipleOuters}.\mathrm{outer}$ yields the $\mathtt{properties}$ of $\mathrm{MyOuter}$ regardless of which route is taken. Because attribute resolution aggregates results via set union (which is an intrinsically idempotent operation), identical property definitions from the two routes merge without conflict. Thus, the calculus eliminates the need for any artificial route-priority or linearization rules.
@@
The rejection occurs because Object1.MyInner and Object2.MyInner each carry the full staticModules of the MyInner deferred module, including the declaration options.outer. When HasMultipleOuters imports both, the module system’s mergeOptionDecls encounters two declarations of the same option and raises a static error; it cannot express that both declarations originate from the same trait and should be unified.
The rejection is analogous to Scala’s “conflicting base types” error above: both systems assume that each option or type member has a single declaration site, and reject the multi-target situation where two inheritance routes introduce the same declaration independently. In inheritance-calculus, the same pattern is well-defined because $\mathtt{overrides}$ (equation 3) recognizes that both routes lead to the same definition, and $\mathtt{supers}$ (equation 2) collects both inheritance-site contexts without duplication.
Appendix E First-Order Semantics of the ANF Fixpoint $\lambda$-Calculus
This appendix reformulates the semantics of Section 3, specialized to $\lambda$-terms via the Single-Path Lemma (Lemma B.1), directly on ANF syntax. The first-iteration approximation gives the lazy $\lambda$-calculus semantics.
The Böhm tree has exactly two node types—scope and application—with separate path types and separate first-order functions for each. All dynamic functions are single-valued: the Single-Path Lemma (Lemma B.1) guarantees that no multi-inheritance arises for translated $\lambda$-terms.